OpenStudy (anonymous):

Given x+y=a, xy=b, express (x-y)^2 in terms of a and b.

OpenStudy (anonymous):

\[\left( x-y \right)^2=x^2+y^2-2xy=x^2+y^2+2xy-2xy-2xy=\left( x+y \right)^2-4xy\] plug the given values and find the solution.

OpenStudy (anonymous):

Thank you very much!